OpenStudy (anonymous):

Please check my answer. I got B. @ganeshie8 @amistre64

OpenStudy (anonymous):

Yup. That's what I got too.

OpenStudy (amistre64):

i would do: 4n*5n(6n)/2, such that 5n = 8.75

OpenStudy (amistre64):

60 n^3 ; n= 8.75/5 60(8.75/5)^3=321.56

OpenStudy (anonymous):

Thank you! :)